• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name gilman

    Log entry time 15:59:47 on March15,2012

    Entry number 366574

    keyword=day shift summary

    8:00 MCC tuning beam
    8:30 Brad says we are going to controlled access, Issue is we cannot get beam back in configuration we were in before. Various chceks and new diagnostics to be done by MCC/acc people.

    Chris Keith has a discussion with Brad, anmd proposes a solution to the polarized target valve problems that we have ben having. It requires that we stay in controlled access all day.

    About 1:20 the Q1s trip. In the middle of haloging it and speaking with Brad and paging Heidi, there is a Hall B fire alarm and we get kicked out of the counting house for about 20 minutes. Heidi works with cryo to reset Q1 while we stand in the parking lot, getting a tan.

    About 2:00: target guys go back into Hall to work on target some more.
    MCC guys continue work on magnets.

    About 3:30: all keys are out to accelerator people and Jack/Vince/Melissa who are doing some DAQ work. Target people cannot get in.

    3:55 techs ramping down Q1s due to low liquid level. Hope to recover soon.